OpenStudy (anonymous):

Someone help? Leroy opened a savings account 25 years ago with a deposit of $3,080.43. The account has an interest rate of 3.7% compounded monthly. How much interest has Leroy earned?

OpenStudy (anonymous):

M=3,080.43(1+(0.037)/(12))^(25*12)

OpenStudy (anonymous):

Hold on .. im going to solve it for you !

OpenStudy (anonymous):

ok cool!!

OpenStudy (anonymous):

I'm so bad with these problems

OpenStudy (ivettef365):

if you go to that website, enter info will give you answer

OpenStudy (anonymous):

I'm doing the same class too -.-

OpenStudy (anonymous):

@ivettef365 is right too ! just go to the site and plug em in but i prefer this one better http://www.thecalculatorsite.com/finance/calculators/compoundinterestcalculator.php

OpenStudy (ivettef365):

which ever you fill more comfortable with, both will give you result

OpenStudy (anonymous):

It didnt give me the right answer

OpenStudy (anonymous):

@Brinley can you help?

OpenStudy (amistre64):

\[B_o(1+\frac r{12})^{12*y}-B_o=interest~only\]

OpenStudy (ivettef365):

what answer did you get ??

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

abt 4677 is what i come up with

OpenStudy (amistre64):

you can also consider this at $1 \[(1+\frac{.037}{12})^{12*25}-1=interest/$\]

OpenStudy (anonymous):

I have some more can you help me?

OpenStudy (amistre64):

1.5183 * 3080.43 = abt 4677 :)

OpenStudy (amistre64):

post them to the left, and im sure someone will bite